A sample of 10 cities with similar populations and monthly sales were assigned different newspaper advertising budgets for one month. The table below summarizes sales​ (in $millions) and newspaper advertising budgets​ (in $thousands) observed during the study.

Sales   Advertising
6.15   3
6.25   3
6.38   6
6.40   6
6.58   9
6.65   9
6.86   12
7.01   12
6.89   15
6.76   15

A. Fit a quadratic regression model and state the quadratic regression.

C. Predict the mean monthly sales for a city with newspaper advertising of

​$17,000.

D. Interpret the meaning of the coefficient of multiple determination.

E. Compute the adjusted r2.
Answer #1

A. Fit a quadratic regression model and state the quadratic regression.

Sales 7.1 7 y = -0.004x2 +0.132x + 5.802 R2 = 0.8944 6.9 6.8 6.7 6.6 5. 5 6.4 6.3 6.2 6.1 0 4 6 8 10 12 14 16 N

C. Predict the mean monthly sales for a city with newspaper advertising of

​$17,000.

> y = 5.802 + 0.132x - 0.004 x2

y = 5.802 + 0.132 * 17 - 0.004 * 17*17

y = 6.89 ($million)

D. Interpret the meaning of the coefficient of multiple determination.

R2 = 0.8944 means that Advertising (K $) explains 89.44% variability in Sales ($Million)

E. Compute the adjusted r2.

R2 = 0.8944

Please rate my answer and comment for doubt.